# Introducing Semantic Capability in LinkedIn’s Content Search Engine

### v1

@@ Abstract.
In the past, most search queries issued to a search engine were short and simple. A keyword based search engine was able to answer such queries quite well. However, members are now developing the habit of issuing long and complex natural language queries. Answering such queries requires evolution of a search engine to have semantic capability. In this paper we present the design of LinkedIn’s new content search engine with semantic capability, and its impact on metrics.
@@ 4. Retrieval layer
We train the two towers in this model using (query, post, label) data collected from historical serving of posts for queries issued to the content search engine. We prepare the label as a combination of the on-topicness and long-dwell labels.
@@ 8. What’s next?
Although the on-topic rate metric does a reasonable job at capturing how a member would perceive the basic quality of posts, its simple definition does not capture all the expectations one would have for the quality of posts for the various types of queries. So, we are evolving the on-topic rate metric to a new metric that would better capture those expectations. Since optimizing this new metric requires a deep understanding of language, we plan to leverage a cross-encoder LLM in the ranking layer. We hope to share our learnings on this in the coming months.

### v2

@@ Abstract.
In the past, most search queries issued to a search engine were short and simple. A keyword based search engine was able to answer such queries quite well. However, members are now developing the habit of issuing long and complex natural language queries. Answering such queries requires evolution of a search engine to have semantic capability. In this paper we present the design of LinkedIn’s new content search engine with semantic capability  [^1] https://www.linkedin.com/search/results/content, and its impact on metrics.
@@ 4. Retrieval layer
We train the two towers in this model using (query, post, label) data collected from historical serving of posts for queries issued to the content search engine. We prepare the label as a combination of the on-topicness and long-dwell labels (label aggregation method).
@@ 8. What’s next?
Although the on-topic rate metric does a reasonable job at capturing how a member would perceive the basic quality of posts, its simple definition does not capture all the expectations one would have for the quality of posts for the various types of queries. So, we are evolving the on-topic rate metric to a new metric that would better capture those expectations. Since optimizing this new metric requires a deep understanding of language, we plan to leverage an LLM in the ranking layer that jointly attends to the query and post text. We hope to share our learnings on this in the coming months.